# ValveFit: An analysis-suitable B-spline-based surface fitting framework for patient-specific modeling of tricuspid valves

### v1

@@ 5 Conclusion
In this article, we introduce ValveFit, a novel GPU-accelerated B-spline surface fitting framework for 3D surface reconstruction of tricuspid heart valve geometry from noisy and sparse point cloud data obtained via medical image segmentation. In this approach, surface fitting is carried out by deforming a template B-spline geometry. We introduce novel regularization terms that maintain the local orthogonality of the surface tangent vectors and constrain the variation of the normal vectors to ensure that the underlying B-spline surface is smooth and intersection-free. The resulting B-spline surfaces are suitable for biomechanical simulations. We first validated the ValveFit framework with a benchmark point cloud dataset that is derived from structural simulation. Here, we also demonstrate robustness toward fitting varying point cloud density and increasing noise intensity. Finally, we have demonstrated surface fitting for real patient data, where we accurately captured the complex three-leaflet structure of the tricuspid valve. The fitted surfaces closely approximate the inherent valve geometry represented by the point clouds across multiple time points in the cardiac cycle. We demonstrate that the fitted surfaces are suitable for biomechanical analysis by performing a structural simulation of the fitted geometry and obtaining strain distribution during valve deformation. These patient-specific valve geometries can offer insight into the progression of valve-related complications, enabling biomechanical modeling-guided precise and effective interventions.
Despite these successes, there are some limitations and scope for improvements for ValveFit. While we were able to generate non-intersecting surfaces without any explicit terms in the loss function to model contact, certain patient-specific scenarios may still lead to leaflet intersections. In our future work, we plan to incorporate an explicit contact term in the loss function to prevent intersections. This will also allow us to fit the closing stages of heart valves more accurately. We also plan to extend the ValveFit framework to support volumetric spline parameterization and patient-specific volumetric modeling for isogeometric analysis.

### v2

@@ 6 Conclusion
In this article, we introduce ValveFit, a novel GPU-accelerated B-spline surface fitting framework for 3D surface reconstruction of tricuspid heart valve geometry from noisy and sparse point cloud data obtained via medical image segmentation. In this approach, surface fitting is carried out by deforming a template B-spline geometry. We introduce novel regularization terms that maintain the local orthogonality of the surface tangent vectors, prevent global self-intersections between surfaces through tangent point energy, and constrain the variation of the normal vectors to ensure that the underlying B-spline surface is smooth and intersection-free. The resulting B-spline surfaces are suitable for biomechanical simulations. We first validated the ValveFit framework with a benchmark point cloud dataset that is derived from structural simulations. Here, we also demonstrate robustness toward fitting varying point cloud densities and increasing noise intensities. Finally, we have demonstrated surface fitting for real patient data, where we accurately captured the complex three-leaflet structure of the tricuspid valve. The fitted surfaces closely approximate the inherent valve geometry represented by the point clouds across multiple time points in the cardiac cycle. We have demonstrated the robustness of the proposed ValveFit framework for point clouds obtained from echocardiographic images of heart valves, which are inherently noisy and have a lower spatial resolution compared to other imaging modalities, such as cine MR or CT. Thus, we envision that the ValveFit framework can be easily extended to point cloud data obtained from other imaging modalities as well. We demonstrate that the fitted surfaces are suitable for biomechanical analysis by performing a structural simulation of the fitted geometry and obtaining the strain distribution during valve deformation. These patient-specific valve geometries can offer insight into the progression of valve-related complications, enabling biomechanical modeling-guided precise and effective interventions.
In our future work, we would like to extend the applicability of the ValveFit framework to patient-specific geometric modeling of the mitral, aortic, and pulmonary heart valves. To apply the proposed framework to these valve types, one simply needs to start from their respective idealized template geometries. Similar to the approach discussed above to construct the template geometry of the tricuspid valve, we can generate anatomical and topologically similar idealized template geometries for the rest of the heart valves, which would serve as the input to the ValveFit framework for patient-specific geometric modeling. We also plan to extend the ValveFit framework to support volumetric spline parameterization and patient-specific volumetric modeling for isogeometric analysis.